OpenStudy (anonymous):

the answer key does not show negatives in my solution but i have two of them. some someone please point out to me my mistakes 3xsquared=5x+4 use the quadratic formula

OpenStudy (anonymous):

I ended up with 

OpenStudy (amistre64):

put everything to one side 3x^2 -5x -4 = 0 then use the quad formula

OpenStudy (anonymous):

I did negative 3x squared instead of the =5x and -4

OpenStudy (anonymous):

is that not okay?

OpenStudy (phi):

you have a negative in the top and bottom, so you can write it as \[ (5 \pm \sqrt{73})/6 \]

OpenStudy (amistre64):

it should come out the same regardless

OpenStudy (amistre64):

the important thing is that you ahve everything to one side before pluging in

OpenStudy (anonymous):

i did that

OpenStudy (amistre64):

notice that you can factor out a negative from top and bottom

OpenStudy (amistre64):

-/- = +

OpenStudy (phi):

so you did it correctly, except for simplifying it a little bit by multiplying top and bottom by -1.

OpenStudy (amistre64):

\[\frac{-5\pm\sqrt{...}}{-6}=\frac{-1}{-1}\frac{5\pm\sqrt{...}}{6}=\frac{5\pm\sqrt{...}}{6}\]

OpenStudy (anonymous):

and that wont effect 73?

OpenStudy (phi):

it changes +/- to -/+ but that is the same thing

OpenStudy (amistre64):

um, no; if anthng the +- goes to -+

OpenStudy (amistre64):

lol .... i gots me a mind reader ;)

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

yep. good luck from all of us